OpenStudy (anonymous):

the tools plus company is planning to produce drills. the company has rented space for its manufacturing operation at $4,000 per month. each drill requires $44 dolars worth of materials and $36 worth of labor. find the total cost if tools plus company makes 330 drills?

OpenStudy (anonymous):

26400?

OpenStudy (anonymous):

seen this one before. each drill cost 44+36=80 to make so compute \[330\times 88+4000\]

OpenStudy (anonymous):

so what would the answer be hun?

OpenStudy (bahrom7893):

Am I a group superhero now?

OpenStudy (anonymous):

i get $30400 but you should use a calculator

OpenStudy (bahrom7893):

YAY I AM!

OpenStudy (anonymous):

33040

OpenStudy (anonymous):

SUPERHERO!. green lantern?

OpenStudy (bahrom7893):

LOL! Good idea, i need to change my prof pic to sum superhero

OpenStudy (anonymous):

thats wrong nvm

OpenStudy (anonymous):

i still get 30400 you should try it for yourself cwc

OpenStudy (anonymous):

Nice Plato

OpenStudy (bahrom7893):

44+36 is not 88 it is 80

OpenStudy (bahrom7893):

330*80 + 4000 =

OpenStudy (bahrom7893):

$30400

OpenStudy (bahrom7893):

I wonder how the heck satellite managed to get the right answer with wrong numbers

OpenStudy (anonymous):

typo. i think i knew that 44+36=80

OpenStudy (anonymous):

typed in in calculator right, in here wrong.